OpenStudy (anonymous):

Find the slope and y-intercept of the line. y = –x + 7 slope: –7, y-intercept: 1 slope: –1, y-intercept: 7 slope: 1, y-intercept: –7 slope: 7, y-intercept: –1

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@abswag13

OpenStudy (anonymous):

ok im here

OpenStudy (anonymous):

slope is -1 intercept is 7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem,

OpenStudy (anonymous):

Marco is selling floral arrangements. Each arrangement uses 1 vase and 11 orchids. Each vase costs Marco $4.00. Let C be the total cost of the arrangement and r be the cost of 1 orchid. Which equation should Marco use to find the total cost of each arrangement? (Points : 1) 11 = C +4r 11C = r + 4 C = 4r + 1 C = 11r + 4

OpenStudy (anonymous):

the correct equation would be the last one ok?

OpenStudy (anonymous):

ok brb other people need me

OpenStudy (anonymous):

ok tyyy

OpenStudy (anonymous):

ok im back

OpenStudy (anonymous):

the correct answer is the last one